at 0°C, the density of mercury is 13.595 x 10³ kg/m³. What is the height of the mercury column in a barometer if the pressure P is Exactly 1 atm = 101.325 kPa?


1
Expert's answer
2021-06-03T05:29:27-0400

Patm = D×g×H

1 = 13.595 × 103 × H

h = 13595/101.325×273

0.4915m